OpenStudy (anonymous):

A total of 120,000 votes were cast for 2 opposing candidates, Garcia and Pérez. If Garcia won by a ratio of 5 to 3, what was the number of votes cast for Pérez?

OpenStudy (sepeario):

Essentially you want to convert the ratios into fractions. If Garcia won by a ratio of 5:3, then he must have won 5/8 of the total candidates. (because 5+3=8)

OpenStudy (sepeario):

Then Perez must have won 3/8 of the candidates. Let's make a variable x to find his number of candidates. \[3/8=x/120000\]

OpenStudy (sepeario):

Can you simplify that and solve for x?

OpenStudy (anonymous):

I'm confused on the changing fractions to ratios

OpenStudy (anonymous):

@minisweet4 Instructions Write the numerator of the fraction. Write a colon (":"). Write the denominator of the fraction. Rewrite the numerator, colon and denominator in simplified terms if necessary. For example, if you have the fraction 5/10, you can convert this to the ratio of 5:10. You can then divide both numbers by 5 to get a simplified ratio of 1:2. The ratio can also be written as "1 to 2." Read more : http://www.ehow.com/how_8430467_convert-fraction-ratio.html

OpenStudy (anonymous):

I have the ratio 5:3 then I change it into the fraction 5/3 then what?

OpenStudy (anonymous):

@MAli13chineta

OpenStudy (anonymous):

im not sure of the rest of the equation sorry-_-